Solve the equation by completing the square. Round to the nearest hundredth if necessary.
x^2-x=1

Answer by lwsshak3(11628)

Solve the equation by completing the square. Round to the nearest hundredth if necessary.
x^2-x=1
(x^2-x+1/4)=1+1/4
(x-1/2)^2=5/4
take sqrt of both sides
x-1/2=±√5/2
x=1/2±√5/2
x=(1+√5)/2
or
x=(1-√5)/2